OpenStudy (theslytherinhelper):

What is halfway between pi and 2pi?

OpenStudy (anonymous):

3pi/2

OpenStudy (anonymous):

pi+pi/2

OpenStudy (mathmale):

Rizwan got off to a good start, but did not finish the calculation of the value halfway between pi and 2pi. He started out finding the mean or average of the two values. To complete his work:\[(\pi+2\pi)/2=\frac{ 3\pi }{ 2 }. \] That's it. Here's another method: imagine you have pi and 2pi plotted on the positive x-axis: pi 2pi Multiply and then divide both of these values by (2/2). You'll get 2pi/2 4pi/2 which are exactly equivalent to pi and 2pi. \[\frac{ 2\pi }{ 2 },\frac{ ? }{ 2 },\frac{ 4\pi }{ 2 }\]

OpenStudy (mathmale):

Bet you can guess what integer to pick to replace that "?"!

OpenStudy (theslytherinhelper):

3?

OpenStudy (mathmale):

Genius!!

OpenStudy (theslytherinhelper):

Thank you!